# Pre-supernova O-C shell mergers could produce more 44Ti than the explosion

### v2

@@ 4 Conclusions
These predicted results do not properly account for the pre-collapse evolution of the $\mathrm{O}$-$\mathrm{C}$ shell merger and how feedback effects may alter the final nucleosynthesis. However, this work does show the importance of including insights from 3D macro physics into 1D stellar evolution and explosion models to better understand the nucleosynthesis of radioactive isotopes in massive stars. Another feature missing from this model is a treatment of rotation as Chieffi & Limongi (2017) found it could boost pre-explosive $\mathrm{{}^{44}Ti}$ production during convective $\mathrm{O}$ shell burning. 3D effects of a convective downturn and boosted mixing speeds compared to MLT are general features of $\mathrm{O}$ shell convection (Jones et al., 2017), and our results find a spread even in the absence of a merger (left column of Fig. 3). 3D explosive effects may also complicate this picture as $\mathrm{{}^{44}Ti}$ production is boosted compared to 1D explosions (Sieverding et al., 2023). We note that both Chieffi & Limongi (2017) and Sieverding et al. (2023) results are included in Fig. 4. The relationship between mergers and the subsequent explosion need further study as the non-radial perturbations from it may also play a significant role in how the subsequent supernova explodes (Couch & Ott, 2013; Müller et al., 2017; Müller, 2020; Andrassy et al., 2020; Yadav et al., 2020; Fields & Couch, 2020; Rizzuti et al., 2024; Laplace et al., 2025).
@@
Acknowledgements.
We would like to thank Marco Pigantari and Roberto Lorenzi for useful discussions and feedback. FH is supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant. This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group at the University of Victoria and hosted on the Digital Research Alliance of Canada Arbutus Cloud at the University of Victoria. These simulations were performed on Digital Research Alliance’s Niagara supercomputer cluster operated by SciNet at the University of Toronto. This work benefited from interactions and workshops co-organized by The Center for Nuclear astrophysics Across Messengers (CeNAM) which is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0023128.
Appendix A $\mathrm{{}^{44}Ti}$ Yields
Here we provide a full list of citations for the yields of $\mathrm{{}^{44}Ti}$ shown in Fig. 4. Table 1 is the predicted pre-explosive $\mathrm{{}^{44}Ti}$ yields from this work that are calculated according to the method described in Issa et al. (2025b). Table 2 provides the pre-explosive and explosive $\mathrm{{}^{44}Ti}$ yields from the massive stars from the NuGrid data set (Pignatari et al. 2016; Ritter et al. 2018) using the delayed explosion mass cut from Fryer et al. (2012). Table 3 is the selection of 1D, 2D, and 3D models from the literature that report explosive $\mathrm{{}^{44}Ti}$ yields. Finally, Table 4 is a list of inferred $\mathrm{{}^{44}Ti}$ yields from observed supernova remnants in the literature.
@@ Appendix A $\mathrm{{}^{44}Ti}$ Yields
Mixing Profile Ingestion Rate ($\mathrm{M}_{\odot}\penalty 10000\ \mathrm{s}^{-1}$) Predicted $\mathrm{{}^{44}Ti}$ Yield ($\mathrm{M}_{\odot}$)
@@
$D_{\mathrm{3D{-}inspired}}$ 0 $1.61\times 10^{-7}$
$D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-5}$ $1.66\times 10^{-7}$
$D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-4}$ $2.46\times 10^{-7}$
$D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-3}$ $1.60\times 10^{-6}$
$3\times D_{\mathrm{3D{-}inspired}}$ 0 $1.85\times 10^{-7}$
$3\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-5}$ $2.22\times 10^{-7}$
$3\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-4}$ $9.54\times 10^{-7}$
$3\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-3}$ $5.65\times 10^{-5}$
$10\times D_{\mathrm{3D{-}inspired}}$ 0 $2.40\times 10^{-7}$
$10\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-5}$ $3.03\times 10^{-7}$
$10\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-4}$ $1.53\times 10^{-6}$
$10\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-3}$ $9.83\times 10^{-4}$
$50\times D_{\mathrm{3D{-}inspired}}$ 0 $2.96\times 10^{-7}$
$50\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-5}$ $3.66\times 10^{-7}$
$50\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-4}$ $1.67\times 10^{-6}$
$50\times D_{\mathrm{3D{-}inspired}}$ $4\times 10^{-3}$ $6.69\times 10^{-3}$
@@
Table 3: Explosive $\mathrm{{}^{44}Ti}$ yields from models in the literature.111
@@
Woosley & Weaver (1995) $1.35\times 10^{-33}{-}2.44\times 10^{-4}$ 1D Models ($11{-}40\;\mathrm{\mathrm{M}_{\odot}}$, $0{-}1\;\mathrm{Z_{\odot}}$)
@@
Sukhbold et al. (2016)aaaaaaYields for this work are from http://doi.org/10.17617/1.b. $(0{-}7.45)\times 10^{-5}$ 1D Neutrino-Driven Supernovae ($9{-}120\;\mathrm{\mathrm{M}_{\odot}}$, $Z_{\odot}$)
Chieffi & Limongi (2017) $(0.89{-}6.0)\times 10^{-5}$ 1D Non-Rotating Models ($13{-}120\;\mathrm{\mathrm{M}_{\odot}}$, $\mathrm{Fe}/\mathrm{H}]=-1,0$)
@@
Table 4: $\mathrm{{}^{44}Ti}$ yields from observations in the literature.222
@@
Borkowski et al. (2010) $2.8^{+2.3}_{-1.7}\times 10^{-5}$ G1.9+0.3 aaaaaaThought to be a Type Ia supernova remnant.
@@
Troja et al. (2014) $(0.3{-}3)\times 10^{-4}$ Tycho’s SNRaaaaaaThought to be a Type Ia supernova remnant.
@@
Wang & Li (2016) $(1.3\pm 0.4)\times 10^{-4}$ Cassiopeia A
@@
$\leq$$3.3\times 10^{-5}$ or $\leq$$2.2\times 10^{-1}$ Vela JuniorbbbbbbUnknown supernova type.
$\leq$$4.8\times 10^{-4}$ Tycho’s SNRaaaaaaThought to be a Type Ia supernova remnant.
$\leq$$3\times 10^{-5}$ G1.9E+0.3aaaaaaThought to be a Type Ia supernova remnant.
$\leq$$4\times 10^{-4}$ Kepler’s SNRaaaaaaThought to be a Type Ia supernova remnant.

### v4

@@ IV Conclusions
These predicted results do not properly account for the pre-collapse evolution of the $\mathrm{O}$-$\mathrm{C}$ shell merger and how feedback effects may alter the final nucleosynthesis. However, this work does show the importance of including insights from 3D macro physics into 1D stellar evolution and explosion models to better understand the nucleosynthesis of radioactive isotopes in massive stars. Another feature missing from this model is a treatment of rotation as Chieffi and Limongi (2017) found it could boost pre-explosive $\mathrm{{}^{44}Ti}$ production during convective $\mathrm{O}$ shell burning. 3D effects of a convective downturn and boosted mixing speeds compared to MLT are general features of $\mathrm{O}$ shell convection (Jones et al., 2017), and our results find a spread even in the absence of a merger (left column of Fig. 3). 3D explosive effects may also complicate this picture as $\mathrm{{}^{44}Ti}$ production is boosted compared to 1D explosions (Sieverding et al., 2023). We note that both Chieffi and Limongi (2017) and Sieverding et al. (2023) results are included in Fig. 4. The relationship between mergers and the subsequent explosion need further study as the non-radial perturbations from it may also play a significant role in how the subsequent supernova explodes (Couch and Ott, 2013; Müller et al., 2017; Müller, 2020; Andrassy et al., 2020; Yadav et al., 2020; Fields and Couch, 2020; Rizzuti et al., 2024; Laplace et al., 2025).
@@
We would like to thank Marco Pignatari, Lorenzo Roberti, and Frank Timmes for useful discussions and feedback. FH is supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant. This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group at the University of Victoria and hosted on the Digital Research Alliance of Canada Arbutus Cloud at the University of Victoria. These simulations were performed on Digital Research Alliance’s Niagara supercomputer cluster operated by SciNet at the University of Toronto. This work benefited from interactions and workshops co-organized by The Center for Nuclear astrophysics Across Messengers (CeNAM) which is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0023128.
Appendix A Conditions of $\mathrm{{}^{44}Ti}$ Production
Fig. 5 provides the temperature and density conditions where $\mathrm{{}^{44}Ti}$ is produced during the $\mathrm{O}$-$\mathrm{C}$ shell merger and how the peak of the nucleosynthesis shifts with different mixing scenarios. As described in §III.1, the higher mixing speeds mix $\mathrm{{}^{44}Ca}$ and $\mathrm{{}^{44}Sc}$ deeper in the shell where temperatures are higher, boosting $\mathrm{{}^{44}Ti}$ production.
Figure 5: The temperature and density conditions for $\mathrm{{}^{44}Ti}$ production. The coloured shading show the net production for $\mathrm{{}^{44}Ti}$ at each mass coordinate for the $D_{\mathrm{MLT}}$ and each of the $D_{\mathrm{3D{-}insp.}}$ mixing profiles with an ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ at $t=110\;\mathrm{\mathrm{s}}$. The solid grey line is the temperature and the dashed grey line is the density for this region. We also include $\rho$-$T$ inset in the lower right corner showing the densities and temperatures where $\mathrm{{}^{44}Ti}$ is produced for all mixing conditions considered in this work.
Appendix B $\mathrm{{}^{44}Ti}$ Yields
Here we provide a full list of citations for the yields of $\mathrm{{}^{44}Ti}$ shown in Fig. 4. Table 1 is the predicted pre-explosive $\mathrm{{}^{44}Ti}$ yields from this work that are calculated according to the method described in Issa et al. (2025b). Table 2 provides the pre-explosive and explosive $\mathrm{{}^{44}Ti}$ yields from the massive stars from the NuGrid data set (Pignatari et al., 2016; Ritter et al., 2018) using the delayed explosion mass cut from Fryer et al. (2012). Table 3 is the selection of 1D, 2D, and 3D models from the literature that report explosive $\mathrm{{}^{44}Ti}$ yields. Finally, Table 4 is a list of inferred $\mathrm{{}^{44}Ti}$ yields from observed supernova remnants in the literature.
@@ Appendix B $\mathrm{{}^{44}Ti}$ Yields
Mixing Profile Ingestion Rate ($\mathrm{M}_{\odot}~\mathrm{s}^{-1}$) Predicted $\mathrm{{}^{44}Ti}$ Yield ($\mathrm{M}_{\odot}$)
@@
$D_{\mathrm{3D{-}insp.}}$ 0 $1.61\times 10^{-7}$
$D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-5}$ $1.66\times 10^{-7}$
$D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-4}$ $2.46\times 10^{-7}$
$D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-3}$ $1.60\times 10^{-6}$
$3\times D_{\mathrm{3D{-}insp.}}$ 0 $1.85\times 10^{-7}$
$3\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-5}$ $2.22\times 10^{-7}$
$3\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-4}$ $9.54\times 10^{-7}$
$3\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-3}$ $5.65\times 10^{-5}$
$10\times D_{\mathrm{3D{-}insp.}}$ 0 $2.40\times 10^{-7}$
$10\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-5}$ $3.03\times 10^{-7}$
$10\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-4}$ $1.53\times 10^{-6}$
$10\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-3}$ $9.83\times 10^{-4}$
$50\times D_{\mathrm{3D{-}insp.}}$ 0 $2.96\times 10^{-7}$
$50\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-5}$ $3.66\times 10^{-7}$
$50\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-4}$ $1.67\times 10^{-6}$
$50\times D_{\mathrm{3D{-}insp.}}$ $4\times 10^{-3}$ $6.69\times 10^{-3}$
@@
Table 3: Explosive $\mathrm{{}^{44}Ti}$ yields from models in the literature.
@@
Woosley and Weaver (1995) $1.35\times 10^{-33}{-}2.44\times 10^{-4}$ 1D Models ($11{-}40\;\mathrm{\mathrm{M}_{\odot}}$, $0{-}1\;\mathrm{Z_{\odot}}$)
@@
Sukhbold et al. (2016)a $(0{-}7.45)\times 10^{-5}$ 1D Neutrino-Driven Supernovae ($9{-}120\;\mathrm{\mathrm{M}_{\odot}}$, $Z_{\odot}$)
Chieffi and Limongi (2017) $(0.89{-}6.0)\times 10^{-5}$ 1D Non-Rotating Models ($13{-}120\;\mathrm{\mathrm{M}_{\odot}}$, $\mathrm{Fe}/\mathrm{H}]=-1,0$)
@@
Table 4: $\mathrm{{}^{44}Ti}$ yields from observations in the literature.
@@
Borkowski et al. (2010) $2.8^{+2.3}_{-1.7}\times 10^{-5}$ G1.9+0.3 a
@@
Troja et al. (2014) $(0.3{-}3)\times 10^{-4}$ Tycho’s SNRa
@@
Wang and Li (2016) $(1.3\pm 0.4)\times 10^{-4}$ Cassiopeia A
@@
$\leq$$3.3\times 10^{-5}$ or $\leq$$2.2\times 10^{-1}$ Vela Juniorb
$\leq$$4.8\times 10^{-4}$ Tycho’s SNRa
$\leq$$3\times 10^{-5}$ G1.9E+0.3a
$\leq$$4\times 10^{-4}$ Kepler’s SNRa